OpenStudy (anonymous):

If 1,000 students take a test that has a mean of 40 minutes, a standard deviation of 8 minutes, and is normally distributed, how many would you expect would finish in between 24 and 48 minutes?

OpenStudy (anonymous):

answer choices: 250 500 750 820

OpenStudy (anonymous):

\[P(24<X<48)=P\left(\frac{24-40}{8}<\frac{X-40}{8}<\frac{48-40}{8}\right)=P(-2<Z<1)\] This will give you the proportion of students that fall in the range of interest. To get the actual number, multiply the proportion by 1000.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw